Profile Productivity Measurement Objective

In 20x2, Choctaw Company implements a new process affecting labor and materials. The following reported data are provided to evaluate the effect on the company’s productivity:

20x1 20x2
Number of units produced 600,000 450,000
Labor hours used 150,000 90,000
Materials used (lbs.) 1,333,333 900,000

Required:

1. Calculate the productivity profile for 20x1. If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x1 Profile
Labor productivity ratio
Material productivity ratio

2. Calculate the productivity profile for 20x2. If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x2 Profile
Labor productivity ratio
Material productivity ratio

3. What if the labor hours used in 20x2 were 127,500? What does comparison of the 20x1 and 20x2 profiles now communicate? If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x1 Profile 20x2 Profile
Labor productivity ratio
Material productivity ratio

Check My Work1 more Check My Work uses remaining.
Answer #1
(1) For 2001
(A) Labor productivity ratio = Units produced / Labor hours used
'= 600,000 / 150,000 = 4 units per labor hour
(B) Material productivity ratio = Units produced / Material used
' = 600,000 / 1,333,333 = 0.45 units per pound
(2) For 2002:
(A) Labor productivity ratio = 450,000 / 90,000 = 5 units per labor hour
(B) Material productivity ratio = 450000 / 900,000 = 0.50 units per pound
Part-3
Labor productivity ratio = 450,000 / 127,500 = 3.53 units per labor hour
Material productivity ratio remains unchanged.